OpenStudy (anonymous):

Identify and draw the following graph: i) 2x^2 + y^2 - 4x + 4y + 6 = 0 Completed the square, and I got: 2(x+1)^2 + (y+2)^2 = 0 I have been taught that you are not suppose to have a number in front of the bracket. With that number there, I am not sure as to how I am suppose to proceed.

OpenStudy (anonymous):

it's a degenerate of an ellipse called a point. ellipse equation: (((x-h)^2)/a)+(((y-k)^2)/b)=r (radius). since radius = 0, it's a point.

OpenStudy (anonymous):

((x+1)^2)+((y+2)^2)/2=0 is ur eqn.

OpenStudy (anonymous):

point in (-1,-2)

OpenStudy (anonymous):

OK I see it, so the "a" would be 1, and the "b" would be sqrt{2}

OpenStudy (anonymous):

no, i just divided both sides by 2.

OpenStudy (anonymous):

it's 1 and 2.

OpenStudy (anonymous):

The Ellipse equation comes in the form: (x - h)^2/a^2 + (y - k)^2/b = 1

OpenStudy (anonymous):

(x - h)^2/a^2 + (y - k)^2/b^2 = 1

OpenStudy (anonymous):

so wouldn't the "b" value be \sqrt{2}?

OpenStudy (anonymous):

oh yes, i'm sorry :)

OpenStudy (anonymous):

How does this actually look like?

OpenStudy (anonymous):

I don't know how to go about with graphing it..

OpenStudy (anonymous):

Still confused about this question. Would someone mind helping me?